# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ Abstract
Quantum coherence has been shown to impact the operational capabilities of quantum systems performing thermodynamic tasks in a significant way, and yet the possibility and conditions for genuine coherence-enhanced thermodynamic operation remain unclear. We show that for steady-state quantum thermal machines —both autonomous and externally driven— that interact weakly with thermal reservoirs and work sources, the presence of coherence induced by perturbations in the machine Hamiltonian guarantees a genuine thermodynamic advantage. Such advantage applies both for the cases in which the induced coherence is between levels with different energies or between degenerate levels. On the other hand, we show that engines subjected to noise-induced coherence, can be outperformed by classical stochastic engines using exactly the same set of (incoherent) resources. We illustrate our results with three prototypical models of heat engines and refrigerators and employ multi-objective optimization techniques to characterize quantum-enhanced regimes in connection with the thermodynamic uncertainty relation and beyond it.
@@ I Introduction
The determination of the interplay between quantum coherence —i.e. the ability of quantum systems to exist in superpositions of multiple states— and thermodynamic operation, constitutes one of the main challenges in quantum thermodynamics, attracting increasing attention during the last decade [1, 2]. One of the main approaches to explore this link, in line with the original spirit of thermodynamics, has consisted in the construction and analysis of minimal models of quantum thermal machines showing different types of coherent evolution [3, 4, 5, 6, 7, 8, 9]. These machines consist of a quantum system —the working medium— composed by few energy levels or qubits which, by coupling to thermal baths at different temperatures and external work sources, are able to perform useful thermodynamic tasks, such as work extraction, heat pumping, or refrigeration. Recent advances in the manipulation and control of quantum systems in the laboratory allowed the implementation of first prototypes, where the basic principles of these models can be mapped to realistic devices in platforms ranging from ion-traps [10, 11] to NV centers in diamond [12], just to mention a few of them [1, 13].
Since the pioneering works of Scully et al. introducing a photo-Carnot engine [14], quantum coherence has been claimed to increase the power output or efficiency of many different types of quantum heat engines and refrigerators. Such improvements are particularly relevant in the case of continuous machines working in steady-state conditions [15], which in principle require less control of the dynamics and couplings with reservoirs. Particularly relevant examples include power-enhancements by noise-induced coherence in lasers, photocell engines, or quantum dots engines [16, 17, 18] (with applications in photosynthetic light harvesting [19]), or by input external coherent fields [9, 12], as well as cooling boosts by degenerate coherence in local models of quantum absorption refrigerators [20, 21, 22]. In all such cases, coherence has been found to play a positive role in the output mechanism, eventually leading to an increased ability of the machine for work extraction or refrigeration. However, since all such output mechanisms are model-dependent, it remains unclear whether the performance shown by these machines cannot be achieved by other equivalent classical models [9, 23, 22], so that a truly quantum thermodynamic advantage can be identified. Given the possibility of implementation of these models in the laboratory and their potential applications, clarifying this point becomes an urgent and crucial point to the field.
@@
In this paper we show how to identify genuine coherence-induced quantum thermodynamic advantage and how to quantitatively characterize it in steady-state quantum thermal machines by combining three main ingredients. The first one is the systematic construction of thermodynamically-equivalent classical thermal machines that are able to produce the same average currents than their quantum-coherent counterparts, while using exactly the same amount of incoherent resources (essentially bath temperatures and energy structure). However, even if same average currents can be reproduced by classical models, fluctuations as captured by the variance of the currents might present significant differences. The comparison of fluctuations in the output currents (reliability of the machine) with respect to the classical counterpart is henceforth the second necessary ingredient in our analysis. This is connected with violations in the so-called thermodynamic uncertainty relation (TUR) [24, 25, 26], which provides a universal trade-off relation between power, efficiency, and output reliability for any classical Markovian thermal machine operating at steady-state conditions [27]. The TUR sets up a model-independent limit in the maximum reliability achievable by any classical machine at a given power output and efficiency. Hence the observation of TUR violations in quantum Markovian machines working in steady-state conditions [28, 29, 30, 31, 32, 33] may be considered as an unambigüous witness of a quantum thermodynamic signature. However the contrary is not necessarily true, i.e. the absence of a TUR violation does not imply the absence of thermodynamic enhancements in the machine, as we explicitly show for different classes of quantum machines, like e.g. autonomous multipartite machines. Moreover, the violation of the TUR in specific parameter regimes is not yet a sufficient (nor necessary) condition for ensuring a practical quantum-thermodynamic advantage in thermal machines, since these parameter regimes may be far from optimal performance and therefore be useless after all. The third ingredient that we employ here is then the use of multi-objective optimization techniques (Pareto optimization) [34, 35, 36, 37, 38, 39] to actually confirm enhanced stability (precision), power output or efficiency, with respect to classical thermodynamically-equivalent machines, in relevant (optimal) regimes of operation of the machine, such as maximum power conditions.
@@
Our general results show systematic quantum-thermodynamic advantages in the presence of Hamiltonian-induced coherence that are applicable to a broad class of steady-state quantum thermal machines (both autonomous or externally driven) in the weak coupling limit. Moreover, we illustrate our results by employing three well-known (prototypical) models showing coherent-induced evolution (see Fig. 1), each corresponding to one of the three possible types of coherence that can arise in the working medium when approaching the steady state: (a) coherence between different energy levels (energetic coherence) induced by external driving, (b) coherence between levels with degenerate energies induced by internal Hamiltonian couplings, and (c) noise-induced coherence on degenerate levels induced by the reservoir. In the case of Hamiltonian-induced coherence (cases a and b) practical quantum-thermodynamic advantages are unambiguously identified, leading to enhanced optimal regimes not allowed by any equivalent classical engine using the same amount of thermodynamic resources. On the contrary, noise-induced coherence between degenerate levels (case c) can lead to disadvantages in the performance, even in cases where the quantum machine dynamics contains intrinsic quantum features.
This paper is organized as follows: Sec. II introduces the different thermal machine models in detail. Sec. III discusses how to characterize the thermodynamic performance in such systems. In Sec. IV, we describe the concept of classical equivalent machines and provide a general recipe for constructing them in the presence of different types of coherence. In Sec. V, we show that in the presence of Hamiltonian coherence, quantum improvements over classical equivalent machines can always be achieved for a broad class of quantum models including unicyclic engines and multipartite (multicycle) engines, whereas in the presence of noise-induced coherence, the performance of quantum machines over their classical counterparts depends on the system parameters. In Sec. VI we verify our general statements using three specific models of quantum thermal machines that use different types of coherence to operate. In Sec. VII we optimize both quantum and classical thermal machines and compare their optimal solutions. Finally, Sec. VIII provides a summary and conclusions of our main results.
II Quantum thermal machines models
@@
We consider thermal machines running continuously in a steady-state regime described as open quantum systems interacting with two or more thermal baths at different temperatures. The machine system has $N$ energy levels, some of which may be degenerate or not, and interconnected through incoherent transitions mediated by the baths. Moreover, we consider the possibility of one or more coherent interactions arising as a consequence of one of the three following sources (see Fig. 1): (a) external driving by a classical field (such as in quantum heat engine models of masers and lasers [3, 4, 5]); (b) internal Hamiltonian interactions between subsystems in few-body machines (e.g. machines composed by several interacting qubits [6] or harmonic oscillators [7]), and (c) noise-induced coherence caused from collective dissipation acting on two or more resonant transitions (as in some light-harvesting complexes [19] and synthetic heat engine models [17]).
@@
$H=H_{0}+V(t),$ (1)
where $H_{0}=\sum_{i=0}^{N-1}\epsilon_{i}\ket{i}\bra{i}$ | is a local Hamiltonian describing $N$ energy levels with $\epsilon_{0}\leq\epsilon_{1}\leq...\leq\epsilon_{N-1}$, and $V(t)$ is a (eventually time-dependent) term capturing coherent transitions between them. It can either represent an external field periodically driving a transition in $H_{0}$, or, in the case of multipartite systems, the internal interaction among machine constituents. An important feature of this interaction Hamiltonian is that it does not severely modify the energy level structure of the system, so that it can be treated as a perturbation to the local Hamiltonian $H_{0}$ (i.e. $|V|\ll|H_{0}|$). On the other hand, for thermal machines using only noise-induced coherence we will typically have $V(t)=0$ and then $H=H_{0}$, since that type of coherence appears solely from the effect of the baths. Without loss of generality, we set to zero the ground-state energy of the machine, $\epsilon_{0}\equiv 0$.
@@
with $\alpha_{ij}^{k}=1$ if $\epsilon_{j}-\epsilon_{i}=\Delta\epsilon_{k}$ and 00 otherwise. In the case of non-degenerate transitions, the above operators reduce to simple jumps $L_{k}^{(r)}=\sqrt{\gamma_{ij}}\ket{j}\bra{i}$ | between energy levels $i\rightarrow j$. However, the operators in Eq. (4) can also describe collective jumps where two or more transitions with same (degenerate) gap $\Delta\epsilon_{k}$ in $H_{0}$ may occur simultaneously, e.g. $i\rightarrow j$ and $i\rightarrow j^{\prime}$ if $\epsilon_{j}=\epsilon_{j}^{\prime}$. In any case, every transition is connected to a single thermal bath, the rates $\gamma_{ij}\geq 0$ being time-independent and verifying the local detailed balance relation $\gamma_{ij}=\gamma_{ji}~{}e^{-\beta_{r}(\epsilon_{j}-\epsilon_{i})}$, with $\beta_{r}=1/{k_{B}T_{r}}$ the inverse temperature of the bath $r$ and $k_{B}$ Boltzmann’s constant. In the long-time limit, the evolution dictated by Eq. (3) converges to a steady-state, verifying $\mathcal{L}(\pi)=0$, where we denote $\pi(t)$ the steady state density operator. We notice that, due to the presence of the time-dependent Hamiltonian $V(t)$, the steady state can show a (residual) periodic time-dependence in the phase in the Schrödinger picture.

### v4

# Certifying quantum enhancements in thermal machines beyond the Thermodynamic Uncertainty Relation

@@ Abstract
Quantum coherence has been shown to impact the operational capabilities of quantum systems performing thermodynamic tasks in a significant way, and yet the possibility and conditions for genuine coherence-enhanced thermodynamic operation remain unclear. Introducing a comparison with classical machines using the same set of thermodynamic resources, we show that for steady-state quantum thermal machines — both autonomous and externally driven — that interact weakly with thermal reservoirs and work sources, the presence of coherence induced by perturbations in the machine Hamiltonian guarantees a genuine thermodynamic advantage under mild conditions. This advantage applies to both cases where the induced coherence is between levels with different energies or between degenerate levels. On the other hand, we show that engines subjected to noise-induced coherence can be outperformed by classical stochastic engines using exactly the same set of (incoherent) resources. We illustrate our results with three prototypical models of heat engines and refrigerators: the three-level amplifier, the three-qubit autonomous refrigerator, and a noise-induced-coherence machine.
@@ I Introduction
The determination of the interplay between quantum coherence —i.e. the ability of quantum systems to exist in superpositions of multiple states— and thermodynamic operation, constitutes one of the main challenges in quantum thermodynamics, attracting increasing attention during the last decade [1, 2]. One of the main approaches to exploring this link, in line with the original spirit of thermodynamics, has consisted of the construction and analysis of minimal models of quantum thermal machines showing different types of coherent evolution [3, 4, 5, 6, 7, 8, 9]. These machines consist of a quantum system —the working medium— composed of few energy levels or qubits which, by coupling to thermal baths at different temperatures and external work sources, are able to perform useful thermodynamic tasks, such as work extraction, heat pumping, or refrigeration. Recent advances in the manipulation and control of quantum systems in the laboratory allowed the implementation of first prototypes, where the basic principles of these models can be mapped to realistic devices on platforms ranging from ion traps [10, 11] to NV centers in diamond [12], just to mention a few of them [1, 13].
Since the pioneering work of Scully et al. introducing a photo-Carnot engine [14], quantum coherence has been claimed to increase the power output or efficiency of many different types of quantum heat engines and refrigerators. Such improvements are particularly relevant in the case of continuous machines working in steady state conditions [15], which in principle require less control of the dynamics and couplings with reservoirs. Particularly relevant examples include power enhancements by noise-induced coherence in lasers, photocell engines, or quantum dots engines [16, 17, 18] (with applications in photosynthetic light harvesting [19]), or by input external coherent fields [9, 12], as well as cooling boosts by degenerate coherence in local models of quantum absorption refrigerators [20, 21, 22]. In all such cases, coherence has been found to play a positive role in the output mechanism, eventually leading to an increased ability of the machine for work extraction or refrigeration. However, since all such output mechanisms are model dependent, it remains unclear whether the performance shown by these machines cannot be achieved by other equivalent classical models [9, 23, 22], so that a truly quantum thermodynamic advantage can be identified. Given the possibility of implementation of these models in the laboratory and their potential applications, clarifying this point becomes an urgent and crucial point for the field.
@@
In this paper, we show how to identify a genuine coherence-induced quantum thermodynamic advantage and how to quantitatively characterize it in steady-state quantum thermal machines by combining two main ingredients. The first one is the systematic construction of thermodynamically equivalent classical thermal machines that are able to produce the same average currents as their quantum-coherent counterparts, while using exactly the same amount of incoherent resources (essentially bath temperatures and energy structure). However, even if the same average currents can be reproduced by classical models, fluctuations, as captured by the variance of the currents, might present significant differences. The comparison of the fluctuations in the output currents (determining the reliability of the machine) with respect to the classical counterpart is henceforth the second necessary ingredient in our analysis. This is connected with violations of the so-called thermodynamic uncertainty relation (TUR) [24, 25, 26], which provides a universal trade-off relation between power, efficiency, and output reliability for any classical Markovian thermal machine operating at steady-state conditions [27]. The TUR sets up a model-independent limit in the maximum reliability achievable by any classical machine at a given power output and efficiency. Hence, the observation of TUR violations in quantum Markovian machines working in steady-state conditions [28, 29, 30, 31, 32, 33] may be considered as an unambiguous witness of a quantum thermodynamic signature. However, the contrary is not necessarily true, i.e. the absence of a TUR violation does not imply the absence of thermodynamic enhancements in the machine, as we explicitly show for different classes of quantum machines, like e.g. autonomous multipartite machines.
@@
Following this method, we certify systematic quantum-thermodynamic advantages in generic steady-state quantum thermal machines (both autonomous or externally driven) that present Hamiltonian-induced coherence in the weak-coupling limit and weak-driving regime. On the other hand, noise-induced coherence can lead to disadvantages on the performance, even in cases where the quantum machine dynamics contains intrinsic quantum features. Moreover, we illustrate our results by employing three well-known (prototypical) models showing coherent-induced evolution (see Fig. 1), each corresponding to one of the three possible types of coherence that can arise in the working medium when approaching the steady state: (a) Hamiltonian-induced coherence between different energy levels (energetic coherence) induced by external driving, (b) Hamiltonian-induced coherence between levels with degenerate energies induced by internal couplings, and (c) noise-induced coherence on degenerated levels induced by the reservoir.
This paper is organized as follows: Sec. II introduces the generic Markovian quantum thermal machine models we use throughout the manuscript. In Sec. III, we describe the concept of classical equivalent machines and provide a general recipe for constructing them in the presence of different types of coherence. In Sec. IV, we present our results for the certification of quantum thermodynamic advantages in the form of two theorems.
In Sec. V we verify our results using three specific examples of quantum thermal machines. Finally, Sec. VI provides a summary and conclusions of our main results. In the appendices, we provide details on the particular models employed to test our results, on the characterization of the thermodynamic performance of quantum thermal machines, and detailed proofs on our main results.
II Markovian quantum thermal machines models
@@
We consider thermal machines running continuously in a steady-state regime, described as open quantum systems that interactct with two or more thermal baths at different temperatures. The machine system has $N$ energy levels, some of which may be degenerate or not, and interconnected through incoherent transitions mediated by the baths. Moreover, we consider the possibility of one or more coherent interactions arising as a consequence of one of the three following sources (see Fig. 1): (a) external driving by a classical field (such as in quantum heat engine models of masers and lasers [3, 4, 5]); (b) internal Hamiltonian interactions between subsystems in few-body machines (e.g. machines composed by several interacting qubits [6] or harmonic oscillators [7]), and (c) noise-induced coherence caused from collective dissipation acting on two or more resonant transitions (as in some light-harvesting complexes [19] and synthetic heat engine models [17]).
@@
$H(t)=H_{0}+V(t),$ (1)
where $H_{0}=\sum_{i=0}^{N-1}\epsilon_{i}\ket{i}\bra{i}$ | is a local Hamiltonian describing $N$ energy levels with $\epsilon_{0}\leq\epsilon_{1}\leq...\leq\epsilon_{N-1}$, and $V(t)$ is a (eventually time-dependent) term capturing coherent transitions between them. It can either represent an external field periodically driving a transition in $H_{0}$, or, in the case of multipartite systems, the internal interaction among machine constituents. An important feature of this interaction Hamiltonian is that it does not severely modify the energy level structure of the system, so that it can be treated as a perturbation to the local Hamiltonian $H_{0}$ (i.e. $|V|\ll|H_{0}|$). On the other hand, for thermal machines using only noise-induced coherence, we will typically have $V(t)=0$ and then $H=H_{0}$, since that type of coherence appears solely from the effect of the baths. Without loss of generality, we set to zero the ground-state energy of the machine, $\epsilon_{0}\equiv 0$.
@@
with $\alpha_{ij}^{k}=1$ if $\epsilon_{j}-\epsilon_{i}=\Delta\epsilon_{k}$ and 00 otherwise. In the case of non-degenerate transitions, the above operators reduce to simple jumps $L_{k}^{(r)}=\sqrt{\gamma_{ij}}\ket{j}\bra{i}$ | between energy levels $i\rightarrow j$. However, the operators in Eq. (4) can also describe collective jumps where two or more transitions with same (degenerate) gap $\Delta\epsilon_{k}$ in $H_{0}$ may occur simultaneously, e.g. $i\rightarrow j$ and $i\rightarrow j^{\prime}$ if $\epsilon_{j}=\epsilon_{j}^{\prime}$. In any case, every transition is connected to a single thermal bath, the rates $\gamma_{ij}\geq 0$ being time independent and verifying the local detailed balance relation $\gamma_{ij}=\gamma_{ji}~{}e^{-\beta_{r}(\epsilon_{j}-\epsilon_{i})}$, with $\beta_{r}=1/{k_{B}T_{r}}$ the inverse temperature of the bath $r$ and $k_{B}$ Boltzmann’s constant (in the following we also take $k_{B}=1$). In the long-time limit, the evolution dictated by Eq. (3) converges to a steady state, verifying $\mathcal{L}(\pi)=0$, where we denote $\pi(t)$ the steady-state density operator. We notice that, due to the presence of the time-dependent Hamiltonian $V(t)$, the steady state can show a (residual) periodic time dependence in the phase in the Schrödinger picture.